OpenStudy (anonymous):

Decide whether x = 2, x = 3 are zeros of f(x) = x^ 4 − x^ 3 − 8x^ 2 + 4x + 12

OpenStudy (anonymous):

Suppose f( 3 ) = 0, then x = 3 are zero of f (x) f (3) = 6 => x = 3 isn't zero! Similarly for x = 2: 2^4 = 16, 2^3 = 8, 2^2 = 4 Can you plug them in?